lolokrat:

can someone help with my geometry homework? i wasn't there for the lecture so im so lost. what i have done is based off of background knowledge and some of them are incomplete. so lost and frustrated.

dude:

Okay lets start with the first triangle (small one) T8gLWTZWS7OaUvF0uXjyng.png

dude:

We know that the angles inside of a triangle must add up to 180 We have 90 (right angle) and 48 (given) 90+48+\(\angle 1\)=180

lolokrat:

<1=42º?

dude:

Ya

lolokrat:

ok how would we do <2 and <3?

lolokrat:

would <2 and <3 be 45º each?

dude:

Not quite, the diagram did not say they were equal

lolokrat:

oh

dude:

Note that there are 2 parallel lines cut by one in between Angle 2 and 39º are alternate interior angles ftiRvXqVQC6ZMX_1RFeLtA.png

dude:

Does that make sense?

lolokrat:

so <3 woul be 39º because they are alt interior

dude:

Right

lolokrat:

<2

dude:

yes angle 2

lolokrat:

that makes sense. would <3 be 51?

dude:

Yes

lolokrat:

ok are we done with that problem?

dude:

Yeah

lolokrat:

is #2 right?

dude:

Yep

lolokrat:

great how do i do #3?

dude:

I assume you know how to do angle 2 6jQ3YEE1RpaygI03zfopQw.png

lolokrat:

your images arent working.

dude:

Oh

dude:

That's weird

1 attachment
dude:

Does that work?

lolokrat:

yup

lolokrat:

would <2 be 29º?

dude:

No, its 2 angles on a straight line They're supplementary lol

lolokrat:

would it be 44º? 180-(90+29+17)=44

dude:

You're overthinking this \(\angle 2\)+ 29 =180

lolokrat:

oh 151?

dude:

Yes

lolokrat:

oh

dude:

For angles 1 and 3, use the rule that all triangles are adding up to 180 \(\angle 3\)+17+151=180 and \(\angle 1\)+90+29=180

dude:

If you have a question on how I got each number, ask

lolokrat:

i understand how you got the numbers. <3=12 <1=61

dude:

Yep

lolokrat:

i understand how to do it once its set up for me. but when i look up a new problem im confused about how to set it up

dude:

Ah, well most of this is closely paying attention to what is given For example for the first problem the indication that the lines \(ED\) and \(FC\) were parallel and therefore you needed to use rules

lolokrat:

that makes sense

dude:

I have to do a bit of homework right now, attempt to do #4 and I'll be back to check on ya

lolokrat:

ok, thanks

lolokrat:

#4) <1+<2+105=180 <1+<2=75 28/2 =37.5 75/2=37.5 <1 and <2 = 37.5º 23+24+<3=180 47=180 <3=133º @dude

dude:

Looks right

lolokrat:

ok im confused on the back

dude:

Can you solve for 1?k6wExGcWSweFcHCniKNnGg.png

dude:

1 attachment
dude:

Right it doesn't load lol